OpenStudy (anonymous):

HELPPPP Isabel is a software saleswoman. Let y represent her total pay (in dollars). Let x represent the number of copies of English is Fun she sells. Suppose that x and y are related by the equation y=+2300+120x

OpenStudy (anonymous):

What is the change in Isabels total pay for each copy of English is fun? What is the total pay if she doesnt sell any of them?

undeadknight26 (undeadknight26):

Lol wow i didnt think you were on here any more! Any who: y= 2300+120x What is the change in Isabels total pay for each copy of English is fun? y= 2300+120x f(x) = 2300 + 120x f(1) = 2300 + 120(1) f(1) = 2420 f(2) = 2300 + 120(2) f(2) = 2540 Understand?

undeadknight26 (undeadknight26):

What is the total pay if she doesnt sell any of them? f(0) = 2300 + 120(0) f(0) = 2300 + 0 f(0) = 2300

undeadknight26 (undeadknight26):

Understand?

OpenStudy (anonymous):

yes thank you!